OpenStudy (anonymous):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21.

OpenStudy (anonymous):

\[f(x)=6x+12-3\] \[f(x)=6x-9\] \[f(x)=y\] \[y=6x-9\] \[f \prime(x) :x=6y-9\] \[\frac{ x+9 }{ 6 }=\frac{ 6y }{ 6 }\] \[y=\frac{ x+9 }{ 6 }\] \[f \prime(x)=\frac{ x+9 }{ 6 }\] \[f \prime(21)=\frac{ 21+9 }{ 6 }=5\]

OpenStudy (anonymous):

Hope i helped you